An aeroplane flies horizontally at 80m/s in still air. If the aviator wishes to fly due south and the wind is blowing from the south-east at 30m/s.
1. What course must he steer.
2. How long will it take him to arrive at his destination 200km away.
1
Expert's answer
2019-03-01T11:43:07-0500

"\overline{v}_{r}" − resulting velocity

"\overline{v}_{a}" − aeroplane velocity

"\overline{v}_{w}" − wind velocity

"v_{rx}", "v_{ax}", "v_{wx}"x-components of the "\overline{v}_{r}", "\overline{v}_{a}" and "\overline{v}_{w}", respectively

"v_{ry}", "v_{ay}", "v_{wy}"y-components of the "\overline{v}_{r}", "\overline{v}_{a}" and "\overline{v}_{w}", respectively

"\lvert\overline{v}_{r}\rvert, \lvert\overline{v}_{a}\rvert, \lvert\overline{v}_{w}\rvert" − absolute values of resulting velocity, aeroplane velocity and wind velocity, respectively


1.

"\overline{v}_{a}+\overline{v}_{w}=\overline{v}_{r}"

"v_{ax}-v_{wx}=0"

"\lvert\overline{v}_{a}\rvert\cdot\sin\beta-\lvert\overline{v}_{w}\rvert\cdot\sin\alpha=0"

"80\cdot\sin\beta-30\cdot\frac{\sqrt{2}}{2}=0"

"sin\beta\approx0.265"

"\beta\approx15.4^\circ"


2.

"S=v\cdot t=\lvert\overline{v}_{r}\rvert\cdot t"

"t=\frac{S}{\lvert\overline{v}_{r}\rvert}"

"\lvert\overline{v}_{r}\rvert=\sqrt{v_{rx}^2+v_{ry}^2}=\sqrt{0+v_{ry}^2}=v_{ry}"

"v_{ay}-v_{wy}=v_{ry}"

"\lvert\overline{v}_{a}\rvert\cdot\cos\beta-\lvert\overline{v}_{w}\rvert\cdot\cos\alpha=v_{ry}"

"\lvert\overline{v}_{r}\rvert=v_{ry}=80\cdot0.964-30\cdot\frac{\sqrt{2}}{2}\approx55.9\;m/s"

"v[km/h]=\frac{v[m/s]\cdot1000}{3600}"

"\lvert\overline{v}_{r}\rvert=\frac{55.9\cdot1000}{3600}\approx15.5\;km/h"

"t=\frac{200}{15.5}\approx12.9\;h"